OpenStudy (anonymous):

ADD.WRITE EACH ANSWER IN LOWEST TERM 4/X+1 + 16/X^2-2X-3

OpenStudy (campbell_st):

\[(4 \times (x -3)/((x+1)(x-3) + 16/((x + 1)(x -3)) = 4x + 4/(x+1)(x -3)\] simplifying further \[4(x + 1)/((x+1)(x-3)) = 4/(x -3)\]

jhonyy9 (jhonyy9):

4 16 4 16 ----------- + ------------- = ------ + ----------------- = (x+1) x^2 -2x -3 (x+1) (x-3)(x+1) 4(x-3) +16 4x-12 +16 4x+4 4(x+1) 4 = ------------- = ------------- = ---------- = ----------- = ----- (x-3)(x+1) (x-3)(x+1) (x-3)(x+1) (x-3)(x+1) x-3